# Vision-driven Autonomous Flight of UAV Along River Using Deep Reinforcement Learning with Dynamic Expert Guidance

### v1

@@ IV METHODOLOGY
Figure 4: System architecture. A human expert collects good trajectory before the training and the transitions are represented by blue arrows. $a^{H}_{t}$ denotes the human expert’s action. The transitions during training are represented by black arrows. $a^{E}_{t}$ denotes the IL expert’s action.
The proposed method is shown in Figure 4 and algorithm 1 to solve the river following task. It requires a demonstration dataset from a human expert, an IL agent to provide guidance, and an RL agent that explores the environment, updates the dataset and learns a control policy.
To train an IL agent to learn from demonstration, a dataset $\mathcal{D}^{E}:\{\tau_{i}\}_{N}$ for $N$ successful trajectories is needed. Each trajectory is a sequence of state-action pairs $\tau={s_{1},a_{1},...s_{t},a_{t}}$. The goal of the expert policy $\pi^{E}$ is to maximize the log-likelihood over $\mathcal{D}^{E}$.
@@
To keep improving the expert policy’s performance during the DRL training, trajectories with episodic rewards larger than the reward threshold $T_{reward}$ will be appended to $\mathcal{D}^{E}$. $T_{reward}$ is $4$ in our experiments. Once the dataset is updated, the expert policy is retrained using the newest $2000$ steps from the dataset until it can mimic at least $95\%$ behaviors from the demonstration.
Algorithm 1 DRL with Dynamic Expert Guidance
Expert collects dataset $\mathcal{D}^{E}$
@@
Initialize policy network $\phi$, value network $\psi$
@@
     if time to update DRL then
         Sample a mini-batch $\mathcal{B}$
         Calculate KL divergence $D_{KL}$
         Calculate loss $\mathcal{L}$
         Update value network $\psi$
         Update policy network $\phi$
@@
     if time to evaluation DRL then
@@
              Update the IL policy $\pi^{E}$
@@
The target of RL is to obtain a policy, defined as a distribution over actions conditioned on states $\pi(a_{t}|s_{t})$, to maximize the long-term discounted cumulative rewards
@@
where $\tau$ is a trajectory and $p_{\pi}(\tau)$ is the distribution of the trajectory under policy $\pi$, and $T$ is the time horizon. Due to Proximal Policy Optimization(PPO)’s scalability and generalization ability, we use PPO [27] as a DRL algorithm to update the parameters of the policy network. In our framework, we use policy loss, value loss, and additional action loss to calculate the total loss and optimize the learning parameters.
The policy loss constrains the biggest possible improvement step on a policy using the data we currently have.
$\mathcal{L}_{policy}=\mathbb{E}\Bigl{[}\mathrm{min}{\Bigl{(}r_{t}(\theta)\hat{A_{t}},clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A_{t}}\Bigr{)}}\Bigr{]}$ (5)
where $r_{t}(\theta)=\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{\theta}_{old}(a_{t}|s_{t})}$ denotes the probability ratio between the current and old policy. $\hat{A_{t}}$ is the advantage. The value function network $V_{\psi}$ is updated by the following loss function:
$\mathcal{L}_{value}=\mathbb{E}\Bigl{[}\Bigl{(}R_{t}-V_{\phi}(s_{t})\Bigr{)}^{2}\Bigr{]}$ (6)
where $R_{t}$ represents the returned value. To encourage the learned policy to be close to the policy learned from the demonstration, we can set a penalty term in the loss function
$\mathcal{L}_{action}=D_{KL}(\pi^{\theta},\pi^{E})=\sum\pi^{\theta}(a_{t}|s_{t})\log\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{E}(a_{t}|s_{t})}$ (7)
where $D_{KL}$ is the KL divergence that measures the divergence between two distributions.
Combining the above loss functions for the DRK with dynamic expert guidance becomes
$\mathcal{L}=c_{policy}\mathcal{L}_{policy}+c_{value}\mathcal{L}_{value}+c_{action}\mathcal{L}_{action}$ (8)
@@ V SIMULATION VERIFICATION
One training and one testing river simulation environment 0(b) are created to train learning algorithms and test their generalization abilities. Specifically, these maps differ in the overall river length, variation of river widths and depths, number of (sharp) turns, bridges, tributaries, and sand islands, whereas the lighting conditions and terrain textures are unified to compare the performance of vision-driven controllers in learning geometrical patterns for the river following task.
Observation images are resized to $128$ x $128$ for both encoding and RL training, and the encoded image is a vector of $1024$ float numbers. Variational Autoencoder [16] is obtained for each map by training 2000 randomly collected images from all maps up to the current training map’s difficulty level. This progressive update of the image encoder depends on the texture coherence across all maps, reduces the impact of the encoding module on RL training, and accounts for the downward compatibility of the trained agent. The encoded mean vector (without Gaussian noise) is used as the observation input to RL algorithms.
We ask an expert to execute the designed river following tasks in the simulator and collect actions and states as expert demonstration data. The UAV is randomly initialized in the valid activity space. The expert tries to correct the pose of the UAV and then controls the UAV to follow the river. Each trajectory is $50$ steps long. We collect $50$ trajectories in the medium environment as $\mathcal{D}^{E}$. The expert policy $\pi^{E}$ is trained by behavior clone using $\mathcal{D}^{E}$. After $50$ epochs of training, the expert policy $\pi^{E}$ has a success rate of $93.27\%$ to predict the expert action from the demonstration and is able to achieve a mean reward of $0.50$ in the training environment. The expert policy is used to generate reference actions during the training of RL agents.
Training is done on a desktop computer with a GeForce 2080 GPU and an Intel Core i7-8700 CPU. Code related to RL and IL is based on StableBaseline3 [28] and Imitation Learning Baseline Implementations [29].
@@ VI-A Training Results
The proposed methods together with baseline methods (BC, PPO, and PPO+StaticBC) are used to train a camera agent in the river following environment. All the methods have been trained for $14000$ steps for $2$ different seeds.
Figure5 shows the training results of different methods in the training environment. The training curve is generated by averaging the training results from different seeds and represented by a solid line of the mean reward with ±$\pm 1$ standard error. The episodic reward is calculated by averaging $100$ past episodes and updated every $1024$ steps. The blue horizon line represents the average of $100$ episodic rewards of the expert policy, which is obtained from 50 episodes of evaluation in $2$ different seeds.
@@
PPO baseline slowly improves throughout the reinforcement learning phase but is still lower than the pre-trained BC baseline. The PPO+DaynamicBC initially returns rewards similar to PPO+StaticBC baseline. Both methods can get a reward at or above the BC after $25000$ steps. However, after the DRL agent’s performance surpasses the BC, PPO+DaynamicBC’s learning curve shows a higher increasing speed and finally reaches a mean episodic reward of $2.0$ at $105000$ steps, while the best mean episodic reward of PPO+StaticBC is less than $1.5$. PPO+StaticBC also shows multiple decrease trends during training, mainly because of the constraint of the static expert policy. PPO+DaynamicBC shows one decrease trend ($110000$ to $13000$) however it recovered to $1.7$ after $10000$ steps. On average, PPO+DaynamicBC’s score is $1.5$ higher than PPO, $0.6$ higher than the BC, and $0.4$ higher than the PPO+StaticBC.
@@
Figure 5: Training results of different approaches in the river following environment. The green horizontal line represents the performance of the pre-trained BC.
VI-B Testing Results
@@
Figure 6: Episode reward distributions of four algorithms on training and testing environments. The horizontal bar represents the mean, the vertical bar represents 25% and 75% percentiles, and the white dot is the median. Statistics from 50 episodes for each algorithm.
Although both algorithms PPO+StaticBC and PPO+DynamicBC can achieve high episodic reward in the training map, the dynamic one still achieves higher mean and median episodic rewards in both maps and leads the static one in terms of the maximum episodic reward by a large margin in the testing map (Figure 6). This verifies the superiority of PPO+DynamicBC algorithm since the IL (BC) model is evolving itself while guiding RL (PPO) model, which acts as a better regularizer compared to the PPO+StaticBC algorithm. PPO baseline and BC algorithms perform the worst in both training and testing maps, due to shortage of runtime expert guidance and limited generalization ability, respectively.
@@
Figure 7: Trajectories of two algorithms PPO+StaticBC and PPO+DynamicBC starting from the same pose in the training map, in two experiments. Star symbols signify the agent position every 50 steps. Trajectory color gets darker as the agent moves further away from the starting position.
@@
To qualitatively compare the navigation efficiency and failure modes of both dynamic and static algorithms, we plot their trajectories in the training map when starting at the same locations (Figure 7). The static algorithm agent can only navigate about $70$% of river length in both experiments, before it failed due to exceeding river spline volume (distracted by the tributary 0(b)) and bumping into the bridge, respectively. This is better visualized by the agent view near blue circles. The dynamic algorithm agent not only finishes whole-river navigation in both experiments but also with fewer steps, which justifies its robustness and efficiency. This can be visualized by the leading checkpoints (star in the figure) of the dynamic algorithm agent, meaning the agent moves a longer effective distance with fewer meaningless movements within the same 50 steps than the static agent. Besides, the dynamic algorithm handles better around river corners with smoother trajectory and larger safe margins to river banks. It also rejects tributary when seeing the main stream ahead by timely heading correction. Both cases can be viewed in figures near the red circles in Figure 7.
@@ VII CONCLUSION
In this paper, we develop two riverine simulation environments to train a UAV agent to autonomously follow the river and avoid obstacles. A novel reinforcement learning method that utilizes the dynamic IL expert has been proposed. The performance of the proposed method has been compared with the baseline methods (PPO, BC, and PPO+StaticBC) in the riverine environments and achieves good qualitative and quantitative results. In the training environment, it achieves at least $0.4$ more rewards than all other methods on average during training and the highest best reward, and it is able to efficiently finish the river course without collision. It has good generalization ability as it also performs best in an unseen and difficult testing environment.
In the near future, tunable environmental light conditions like light directions and intensities can be enabled in the simulated riverine gym environment to increase policy robustness in the way of domain randomization and transfer the learned river following ability to a real UAV platform and test it in real-world riverine environment. On the other hand, we will systematically study and analyze the proposed reinforcement learning framework, keep optimizing the performance of the method, and apply it to solve more robotic challenges.

### v2

@@ IV METHODOLOGY
The proposed method, illustrated in Fig. 1 and detailed in Algorithm 1 for solving the river following task, relies on several components: a demonstration dataset provided by a human expert, a BC expert offering guidance, and a PPO agent responsible for environment exploration, dataset updates, and policy learning. The BC expert is firstly trained on the demonstration dataset, $\mathcal{D}^{E}:\{\tau_{i}\}_{N}$ for $N$ successful trajectories, resulting in the final form of StaticBC and initial form of DynamicBC baselines. Each trajectory is a sequence of state-action pairs $\tau={s_{1},a_{1},...,s_{t},a_{t}}$. The goal of the expert policy $\pi^{E}$ is to maximize the log-likelihood over $\mathcal{D}^{E}$.
@@
To enhance the performance of the expert policy during PPO training, trajectories with episodic rewards exceeding the reward threshold $T_{reward}$ are added to dataset $\mathcal{D}^{E}$. $T_{reward}$ is set to $4$ in the riverine environment, $18$ in the CiffCircular environment. Evaluation of the PPO agent occurs every $10000$ training steps in the CliffCircular environment and every $5000$ steps in the riverine environment, with more frequent updates to the expert policy in the latter due to its higher complexity. In the simpler CliffCircular environment with multi-discrete observation, duplicate transitions are removed from dataset $\mathcal{D}^{E}$ to mitigate bias. The expert policy undergoes retraining using the latest $2000$ steps from the dataset for $20$ epochs.
Algorithm 1 Synergistic Reinforcement and Imitation Learning
Collect the transition dataset $\mathcal{D}^{E}$ by the human expert
@@
Initialize policy network $\phi$, value network $\psi$ of RL agent
@@
     if time to update RL agent then
         Sample a mini-batch $\mathcal{B}$ from $\mathcal{D}^{R}$
         Generate actions $a_{t}^{E}$ using $\pi^{E}$
         Calculate the cross entropy $H(\pi^{E},\pi^{\psi})$ (Eq. 4)
         Calculate loss $\mathcal{L}$ (Eq. 5)
         Update value network $\psi$ and policy network $\phi$
@@
     if time to evaluate RL agent then
@@
              Remove repeated transitions from the $\mathcal{D}^{E}$
              Update the IL policy $\pi^{E}$ (Eq. 2 )
@@
The target of PPO is to obtain a policy, defined as a distribution over actions conditioned on states $\pi(a_{t}|s_{t})$, to maximize the long-term discounted cumulative rewards
@@
where $\tau$ is a trajectory and $p_{\pi}(\tau)$ is the distribution of the trajectory under policy $\pi$, and $T$ is the time horizon.
To encourage the PPO policy to be close to the policy learned from the decent demonstrations, we add a penalty term in the loss function
$\mathcal{L}_{action}=H(\pi^{E},\pi^{\psi})=-\sum\pi^{E}(a_{t}|s_{t})\log\pi^{\psi}(a_{t}|s_{t})$ (4)
where $H(\pi^{E},\pi^{\psi})$ is the cross entropy that measures the difference between the expert policy $\pi^{E}$ and the RL policy $\pi^{\psi}$.
Combining the above loss functions, PPO+DynamicBC becomes
$\mathcal{L}=w_{1}\mathcal{L}_{policy}+w_{2}\mathcal{L}_{value}+w_{3}\mathcal{L}_{action}$ (5)
where we keep the policy loss $\mathcal{L}_{policy}$ and value loss $\mathcal{L}_{value}$ of PPO intact. $w_{1}$, $w_{2}$, and $w_{3}$ represent the weight constants associated with each loss. $w_{1}$ and $w_{2}$ are set to 1. $w_{3}$ is set to $1$ at the beginning. If the PPO achieves a mean score better than BC during evaluation, $w_{3}$ is set to $0.2$ to encourage exploration. PPO is updated every $1024$ training steps.
@@ V SIMULATION VERIFICATION
In the riverine environment, observation images are resized to $128$ x $128$ for both encoding and RL training, and the encoded image is a vector of $1024$ float numbers. VAE model is obtained by training $2000$ randomly collected images from the environment. The encoded mean vector (without Gaussian noise) is used as the observation input to RL algorithms. $50$ trajectories each with $50$ steps are collected by a human expert to form the dataset $\mathcal{D}^{E}$. After $50$ epochs of training, the expert policy $\pi^{E}$ has a success rate of $93.27\%$ to predict the expert action from the demonstration and is able to achieve a mean reward of $0.60$ in the riverine environment. This expert policy will be used to generate reference actions during the training of RL agent.
In the CliffCircular environment, $100$ trajectories are collected by an expert, which achieves a mean episode reward of approximately 00. The BC expert, pre-trained on this dataset, attains a reward of $-0.24$, indicating a suboptimal policy that is capable of accomplishing the track following objective but commits few errors occasionally. Code related to RL and IL is based on StableBaseline3 [34] and Imitation Learning Baseline Implementations [35].
@@ VI-A Training Results
The proposed method (PPO+DynamicBC), together with baseline methods (StaticBC, DynamicBC, PPO, and PPO+StaticBC), are used to train the discrete-action agents in both environments. All the methods have been trained for $14000$ steps in the riverine environment and $10000$ steps in the CliffCircular environment for $5$ different seeds.
Fig. LABEL:fig:training_result illustrates the training results of different methods in two environments. The training curve is derived by averaging results from $5$ different seeds and is depicted as a solid line representing the mean reward, with a shaded region around it indicating the standard error. The episodic reward is computed by averaging over $100$ past episodes, with updates occurring every $1024$ steps. Additionally, the green horizon line signifies the performance achieved by a pre-trained (BC) model.
In the CliffCircular environment, the PPO baseline slowly improves and surpasses the pre-trained BC baseline after $85000$ steps. The PPO+DynamicBC initially returns rewards similar to PPO+StaticBC baseline. However, PPO+DynamicBC’s learning curve shows a higher increasing speed after $10000$ steps and finally reaches a mean episodic reward of $17.90$ at $85000$ steps, while the maximum possible reward is $20$. The best mean episodic reward of PPO+StaticBC is less than $11.31$. After $10000$ steps of training, PPO+DynamicBC’s mean episode length is $28.50$, while PPO+StaticBC and PPO’s mean episode lengths are $32.62$ and $88.40$, respectively. This suggests that the proposed method has acquired a more efficient policy for completing the task in fewer steps, whereas PPO+StaticBC has developed a less optimal policy, requiring an additional $4.12$ steps to complete the task. Evidently, PPO encounters challenges in learning an effective policy within this POMDP track following environment with non-Markovian rewards.
@@
In the river following environment, PPO baseline slowly improves throughout the reinforcement learning phase but is still lower than the pre-trained BC baseline. The PPO+DynamicBC initially returns rewards similar to PPO+StaticBC baseline. Both methods can get a reward at or above the BC after $25000$ steps. However, after the PPO agent’s performance surpasses the BC, PPO+DynamicBC’s learning curve shows a higher increasing speed and finally reaches a mean episodic reward of $2.0$ at $105000$ steps, while the best mean episodic reward of PPO+StaticBC is less than $1.5$. PPO+StaticBC also shows multiple decrease trends during training, mainly because of the constraint of the static expert policy. PPO+DynamicBC shows one decrease trend ($110000$ to $13000$) however it recovered to $1.7$ after $10000$ steps. On average, PPO+DynamicBC’s score is $1.5$ higher than PPO, $0.6$ higher than BC, and $0.4$ higher than PPO+StaticBC.
@@
Figure 5: Training results of different approaches in the CliffCircular environment and river following environment.
VI-B Evaluation Results
@@
Figure 6: Episode reward distributions of four algorithms on CliffCircular-gym-v0 and riverine environments. The horizontal bar represents the mean, the vertical bar represents 25% and 75% percentiles, and the white dot is the median. Statistics from 50 episodes for each algorithm.
As shown Fig. 6, the reward and navigation efficiency are compared through violin plots. To generate the episodic reward distribution, all methods with the best weights are evaluated for $50$ epochs in the same seed. In the CliffCircular environment, BC has both high and low scores, resulting in a relatively low mean reward. PPO without guidance learns how to avoid collisions but cannot figure out the task objective. PPO+StaticBC is able to increase the chance of finishing the task and decrease the chance of collision compared to StaticBC. PPO+DynamicBC achieves the best result. It learns an optimal policy to finish the track following task while avoiding cliffs. In the riverine environment, BC and the vanilla PPO can never achieve any reward larger than $6$ while both algorithms PPO+StaticBC and PPO+DynamicBC can achieve higher episodic reward of $10$. Noticeably, PPO+DynamicBC achieves the highest mean and median episodic rewards compared to all other methods.
This verifies the superiority of PPO+DynamicBC algorithm since the IL (BC) model is evolving itself while guiding RL (PPO) model, which acts as a better regularizer compared to PPO+StaticBC algorithm. Meanwhile, PPO and BC baseline algorithms have inferior performance due to the shortage of runtime expert guidance and limited generalization ability, respectively. The difference in training convergence and evaluation performance further illustrates the difficulty of the track following tasks. More importantly, this synergistic combination of PPO and BC is validated to largely boost the training speed and final model performance.
@@
Figure 7: Trajectories of two algorithms PPO+StaticBC and PPO+DynamicBC starting from the same pose in the Unity riverine environment, in two evaluation experiments. Star symbols signify the agent’s position every 50 steps. Trajectory color gets darker as the agent moves further away from the starting position.
@@
To qualitatively compare the navigation efficiency and failure modes of both dynamic and static algorithms, we plot their trajectories in the riverine environment when starting at the same locations (Fig. 7). The static algorithm agent can only navigate about $70$% of river segments in both experiments, before it failed due to exceeding river spline volume (distracted by the tributary 2(b)) and bumping into the bridge, respectively. This is better visualized by the agent view near blue circles. The dynamic algorithm agent not only finishes whole-river navigation in both experiments but also with fewer steps, which justifies its robustness and efficiency. This can be visualized by the leading checkpoints (star in the figure) of the dynamic algorithm agent, meaning the agent moves a longer effective distance with fewer meaningless movements within the same $50$ steps than the static agent. Besides, the dynamic algorithm handles better around river corners with smoother trajectory and larger safe margins to river banks. It also rejects tributary when seeing the main stream ahead by timely heading correction. Both cases can be viewed in figures near the red circles in Fig. 7.
@@ VII CONCLUSION
In this paper, we develop a photo-realistic riverine simulation environment to train a UAV agent to autonomously follow the river while avoiding obstacles. A novel framework that synergistically combines RL and IL is proposed, and evaluated in both a riverine environment and a customized grid world track following environment. The performance of the proposed method (PPO+DynamicBC), compared with the baseline methods (PPO, StaticBC, DynamicBC and PPO+StaticBC) in both environments, achieves the best qualitative and quantitative results. In the riverine environment, it achieves at least $0.4$ more rewards than all other methods on average, and the highest best and median rewards. In the CliffCircular environment, this algorithm finishes the track following task with nearly full score (episode reward $19.04$ out of $20$), with significantly lower reward variance than all baseline algorithms.
Further research can improve the proposed method in several aspects. Firstly, instead of relying solely on the episodic reward threshold to filter trajectories, specific transitions can also be evaluated by some advantage measure and the valuable ones can be merged to the dataset for IL expert refinement. This could reduce the demonstration buffer size, bias of this dataset, and the overall training time. Secondly, IL expert’s policy might not consistently outperform that of the RL agent during training process. Therefore, when to utilize the action loss for updating the RL agent could be further investigated. Thirdly, observation augmentation (with historical observations or actions) techniques can be implemented to push the partially observable non-Markovian decision process to MDP for convergence guarantee in the track following tasks.